ANALYSIS OF THE ROTATIONAL STRUCTURE IN THE HIGH-RESOLUTION INFRARED SPECTRUM OF TRANS-HEXATRIENE-1-^{13}C_{1}

\underline{NORMAN C. CRAIG} and HENGFENG TIAN, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; THOMAS A. BLAKE, Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.

Hexatriene-1-^{13}C_{1} was synthesized by reaction of 2,4-pentadienal and (methyl-^{13}C)-triphenylphosphonium iodide (Wittig reagent). The trans isomer was isolated by preparative gas chromatography, and the high-resolution (0.0015 cm^{−1}) infrared spectrum was recorded on a Bruker IFS 125HR instrument. The rotational structure in two C-type bands was analyzed. For this species the bands at 1010.7 and 893.740 cm^{−1} yielded composite ground state rotational constants of A_{0} = 0.872820(1), B_{0} = 0.0435868(4), and C_{0} = 0.0415314(2) cm^{−1}. The ground state rotational constants for the 1-^{13}C species were also predicted with Gaussian 03 software and the B3LYP/cc-pVTZ model. After scaling by the ratio of the observed and predicted ground state rotational constants for the normal species, the predicted ground state rotational constants for the 1-^{13}C species agreed within 0.005 % with the observed values. Similar good agreement between observed and calculated values (0.016 %) was found for the three ^{13}C species of the cis isomer.^{a} We conclude that ground state rotational constants for single heavy atom substitution can be calculated with adequate accuracy for use in determining semi-experimental equilibrium structures of small molecules. It will be unnecessary to synthesize the other two ^{13}C species of trans-hexatriene.

aR. D. Suenram, B. H. Pate, A. Lesarri, J. L. Neill, S. Shipman, R. A. Holmes, M. C. Leyden, N. C. Craig J. Phys. Chem. A 113, 1864-1868 (2009).